# A continuous boostlet transform for acoustic waves in space–time

### v1

@@ Abstract
Sparse representation systems that encode the signal architecture have had an exceptional impact on sampling and compression paradigms. Remarkable examples are multi-scale directional systems, which, similar to our vision system, encode the underlying architecture of natural images with sparse features. Inspired by this philosophy, the present study introduces a representation system for acoustic waves in 2D space–time, referred to as the boostlet transform, which encodes sparse features of natural acoustic fields with the Poincaré group and isotropic dilations. Continuous boostlets, $\psi_{a,\theta,\tau}(\varsigma)=a^{-1}\psi\left(D_{a}^{-1}B_{\theta}^{-1}(\varsigma-\tau)\right)\in L^{2}(\mathbb{R}^{2})$, are spatiotemporal functions parametrized with dilations $a>0$, Lorentz boosts $\theta\in\mathbb{R}$, and translations $\smash{\tau\in\mathbb{R}^{2}}$ in space–time. The admissibility condition requires that boostlets are supported away from the acoustic radiation cone, i.e., have phase velocities other than the speed of sound, resulting in a peculiar scaling function. The continuous boostlet transform is an isometry for $L^{2}(\mathbb{R}^{2})$, and a sparsity analysis with experimentally measured fields indicates that boostlet coefficients decay faster than wavelets, curvelets, wave atoms, and shearlets. The uncertainty principles and minimizers associated with the boostlet transform are derived and interpreted physically.
@@ 1 Introduction
The analysis and processing of acoustic fields has seen numerous developments since the first computer simulations in the 1960s [3, 4], when applications of geometrical-acoustics methods, e.g., ray tracing [5], quickly became popular. Some years later, the image source method was developed as an analytical tool to predict acoustical reflections in complex room geometries [6]. In the 1980s, Williams and collaborators made monumental advancements in sound source quantification by introducing acoustic holography [7], an industrial breakthrough for contactless characterization of vibroacoustic sources. By the end of the 1990s, Berkhout et al. studied the analysis [8] and extrapolation [9] of acoustic fields inside rooms in great detail by using analogies from seismic exploration. Numerical approaches such as finite elements [10], finite-difference-time-domain methods [11], spectral element methods [12], and discontinuous Galerkin methods [13] have since been used extensively to approximate the solution by discretizing the computational domain. With the emergence of compressed sensing in the mid-2000s [14], numerous studies have investigated the sparse reconstruction of acoustic fields with monopole sources [15, 16], plane waves [17], curvelets [18], and shearlets [19]. Finally, data-driven methods, such as physics-informed neural networks [20, 21, 22, 23, 24], deep prior approaches [25, 26, 27], generative models [28, 29, 30, 31], and neural operators [32, 33, 34, 35], have recently achieved impressive results due to their ability to more closely adapt to the structure (spatial or temporal) of the acoustic fields.
1.1 Sparse representation machinery
Sparse representations have had impactful applications in imaging sciences, such as compressed sensing, super-resolution, multi-scale analysis, and feature extraction for model-based and machine-learning frameworks. Indeed, sampling and reconstructing acoustic fields in space–time is challenging because many data points are required to avoid aliasing. For example, in the human audible range ($20$ Hz to $20$ kHz), about $1.6$ million microphone positions are necessary to reconstruct the pressure field over a cubic meter of air. Thus, overcoming sampling limitations has a profound impact on technologies based on the wave equation, such as seismic exploration, medical ultrasound, and vibroacoustic imaging, to mention a few.
@@ 1.1 Sparse representation machinery
Decades of efforts in applied harmonic analysis have led to remarkable representation systems. In 1946, Gabor presented a theory for time–frequency analysis [36] by applying Heisenberg’s uncertainty principle and establishing an analogy between sound and quanta. Gabor envisioned that a family of parametrized atoms could detect transient components more effectively than Fourier bases. Gabor’s ideas later triggered the development of the wavelet transform in the 1970s, coined initially as the cochlear transform by Zweig [37], and later formalized in various ways by Grossmann, Morlet, Debauchies, and Meyer in the 1980s-1990s [38, 39, 40]. An interesting property of Gabor wavelets is their optimal uncertainty in time and frequency domains simultaneously. Wavelets sit at the core of JPEG2000 compression and offer optimally sparse representations of point singularities in high-dimensional spaces.
The beauty of wavelets lies in their transcendental application in many fields of science and technology. Their limitation, however, is the representation of extended singularities, because of which various anisotropic transforms were constructed in the late 1990s / early 2000s, e.g., ridgelets to represent ridge functions [41], and curvelets [42], shearlets [43], and contourlets [44] to represent curved singularities. Similar variants have been later proposed, such as hyperbolic shearlets [45], alpha molecules [46] and bendlets [47]. These representation systems have been applied successfully to image denoising [48, 49], edge detection [50, 51], and inpainting [52, 53]. Research in microlocal analysis shows that curvelets and shearlets resolve the wavefront set [54, 55]. In the context of acoustics, curvelets [56] and wave atoms [57] have been shown to provide optimally sparse representations of the wave propagator in free space. In this sense, they are functions defined on a so-called space–time foliation, e.g., wave fields evolving from one instant to another with Hamiltonian flows. Demanet and Ying [57] have shown that optimally sparse representations of Fourier integral operators and Green’s function of the wave equation are achieved with anisotropic essential support $\sim 2^{-aj}\times 2^{-j/2}$, where $a\in[1/2,1]$. Despite the remarkable progress, literature is scarce on representations in full space–time [58, 59, 60, 61].
1.2 Multi-scale directional analysis in acoustic scattering and beyond
Modeling localized wave phenomena in the near-field, for example, the scattered wave highlighted in Fig. 1(b), is challenging when the size of the scattering object is comparable to the acoustic wavelength $\lambda$. From the viewpoint of representation sparsity of stationary plane waves, many coefficients must interfere (constructively and destructively) to replicate the wave decay in space–time—especially in the vicinity of the object. Acoustic scattering problems are typically boundary value problems where a sound source ensonifies some finite or infinite body, i.e. causes the latter to scatter the sound wave. The classical approach is to view the total sound field as a sum of the ensonifying sound wave in a completely free field, $p_{\text{ff}}(\mathbf{r},t)$, and a scattered sound wave, $p_{\text{sc}}(\mathbf{r},t)$, which is also assumed to radiate in free space,
$p(\mathbf{r},t)=p_{\text{ff}}(\mathbf{r},t)+p_{\text{sc}}(\mathbf{r},t),$
where $\mathbf{r}=(x,y,z)\in\mathbb{R}^{3}$ and $t\in\mathbb{R}$ are the coordinates of 4D space–time. The sum of these two sound field components must fulfill the specified boundary condition in the free space (that is, a non-reflective boundary condition) at the location of the scattering body’s surface. An orthogonal set of known mode functions (parameterized solutions) exists for a few geometries, such as a sphere, a cylinder, etc. [62]. For these geometries, the boundary condition can be fulfilled mode by mode, and the mode amplitudes for $p_{\text{sc}}(\mathbf{r},t)$ can be derived. Numerical methods based on the Kirchhoff–Helmholtz integral equation can also predict $p_{\text{sc}}(\mathbf{r},t)$. The free-field function $p_{\text{ff}}(\mathbf{r},t)$ is smooth since it is a propagating sound wave. Therefore, $p_{\text{sc}}(\mathbf{r},t)$ must also be a smooth function. Nonetheless, corners and edges of the scattering body may cause significant gradients in the pressure fields.
Taking scattering as a challenging applied problem, this study introduces a sparse representation system for localized waves in space–time and wavenumber-frequency domains, preliminarily investigated by Zea and Laudato in [60]. Experimental evidence from data in that study is shown in Fig. 1. In particular, a scattered wave is shown in the zoomed-in box of Fig. 1(b), which has a space–time decay due to the finiteness of the table in Fig. 1(a). Specifically, we see a wave that is attenuated closer to the left end of the array but which is amplified closer to the center. Additionally, the Fourier spectrum in Fig. 1(c) exhibits hyperbolae localized inside a phase velocity cone. These hyperbolae lie in the co-direction of the waves in space–time and correspond to their phase velocities. Here, we mention acoustic scattering as an application for boostlets, and other applications (see above) may further convince the reader of the potential impact. We envision opportunities that benefit technologies based on the wave equation and potentially other partial differential equations.
Figure 1: Illustration of a room-acoustical measurement in 2D space–time. (a) Experimental setup with a sound source emitting a frequency sweep between $100$ Hz and $4.5$ kHz and a free-field microphone recording the sound pressure. (b) Sound pressure field measured for approximately $8.9$ ms on a line of $100$ microphone positions spanning $3$ m. Zoom-in shows a wave decaying in space–time, entailing a scattered wave. (c) Fourier spectrum of (b). The spectrum shows several hyperbolae localized in a phase velocity cone (dashed lines shown for illustration purposes).
@@ 1.3 Contribution of the work
Previous applications of multi-scale directional transforms to interpolate wavefronts in space–time have proved empirically successful [18, 19]. As mentioned above, these transforms were rightfully designed for image processing, that is, for extended singularities in real space, but not explicitly for waves in full space–time. The central distinction with the boostlet transform is that it considers a unified treatment of space–time for wave propagation problems; specifically, boostlets encode the physical attributes of a localized wave in a manner equivariant to boosts (or hyperbolic rotations), dilations, and translations in space–time. We find the application of geometric tools from special relativity to sparsely represent waves genuinely fascinating and by no means evident. To the best of our knowledge, the application of such special relativity ideas to represent acoustic waves is new.
@@
    We apply the principle of analogy from sparsely coding natural images [63, 64] to sparsely coding acoustic waves in space–time. Previous theoretical and experimental results [60] indicate that natural acoustic fields in 2D space–time can be sparsely encoded into band-limited wave shapes that are neither planar nor spherical. As Fig. 1(c) suggests, anti-symmetric hyperbolae localized in phase velocity cones characterize these shapes. In this work, we will see that this particular structure can be encoded using a certain group structure (that of the Poincaré group).
@@
    We introduce the continuous boostlet transform, which decomposes acoustic waves into a collection of band-limited functions on $L^{2}(\mathbb{R}^{2})$, referred to as boostlets, parametrized with the Poincaré group (Lorentz boosts and translations), and isotropic dilations in 2D space–time. The physical meaning of the admissibility condition of the transform and its peculiar scaling function are discussed. When applied to acoustic fields measured experimentally, boostlets exhibit similarities with scattered waves due to their spatiotemporal decay and appear to have a faster coefficient decay compared to benchmark systems such as wavelets, shearlets, and wave atoms.
  • (iii)
    We derive the uncertainty principles associated with the continuous boostlet transform and the corresponding minimizers for the transform parameter pairs. We discuss their physical interpretation and compare them with the uncertainty principles of 2D wavelets [65] and 2D shearlets [66]. The infinitesimal generators of the dilation and boost operators commute, indicating the existence of a continuum of wave shapes with various Fourier support sizes that are all optimally concentrated with respect to dilation and boosting, including the canonical cases of planar and spherical waves.
@@ 2 Background
Although solutions of the wave equations have a wide frequency range, this study focuses on the acoustic regime between infrasound and ultrasound. Our ambition is that the methodology can be extended to other types of waves and propagation media. This section presents the relevant background on canonical solutions to time-harmonic wave equations, Fourier acoustics theory, and encoding dispersion relations with a group structure.
@@ 2.1 Wave equation preliminaries
[…]
@@
$\nabla^{2}=\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}+\frac{\partial^{2}}{\partial z^{2}}.$
@@
The spatial Fourier transform of Eq. (4) gives the following Dirac delta located at $(\hat{k}_{x},\hat{k}_{y},\hat{k}_{z})$ [67]
$\hat{p}(\mathbf{k};\omega_{0})=\int_{\mathbb{R}^{3}}P(\omega_{0})e^{i\,(\mathbf{r}\cdot\hat{\mathbf{k}})}e^{-i(\mathbf{r}\cdot\mathbf{k})}\mathrm{d}\mathbf{r}=P(\omega_{0})\delta(k_{x}-\hat{k}_{x},k_{y}-\hat{k}_{y},k_{z}-\hat{k}_{z}),$
@@
Experimental evidence [17] suggests that plane wave expansions can sparsely represent stationary acoustic fields in rooms at low frequencies (i.e., the modal region). In this frequency region, it is known that the acoustic field in the room is composed of a few modes, which can be represented by a few plane waves. Similar reports have indicated that monopole sources provide sparse representations of compact vibroacoustic sources [68, 69], which are fully localized in real space and broadband in the wavenumber domain.
@@ 2.2 Acoustic radiation in Fourier space
$\hat{p}(\mathbf{r},\omega_{0})=P(\omega_{0})e^{i(k_{x}x+k_{y}y+k_{z}z)}$
@@
$\hat{p}(\mathbf{r},\omega_{0})=P(\omega_{0})e^{i(k_{x}x+k_{y}y)}e^{-k_{z}^{\prime}z},$
@@
For all values of $(k_{x},k_{y})\in\mathbb{R}^{2}$ and a specific frequency $\omega_{0}$, propagating and evanescent waves are thus divided by the so-called radiation circle [67]. As illustrated in Fig. 2, the extrusion of the radiation circle as the frequency $\omega\in\mathbb{R}$ varies results in the radiation cone. The presence of this cone suggests that a natural decomposition of the wavefield must preserve this cone.

### v2

# A boostlet transform for wave-based acoustic signal processing in space–time

@@ Abstract
Sparse representation systems that encode signal architecture have had a profound impact on sampling and compression paradigms. Remarkable examples are multi-scale directional systems, which, similar to our vision system, encode the underlying architecture of natural images with sparse features. Inspired by this philosophy, we introduce a representation system for wave-based acoustic signal processing in 2D space–time, referred to as the boostlet transform, which encodes sparse features of natural acoustic fields using the Poincaré group and isotropic dilations. Boostlets are spatiotemporal functions parametrized with dilations, Lorentz boosts, and translations in space–time. Physically speaking, boostlets are supported away from the acoustic radiation cone, i.e., having broadband frequency with phase velocities other than the speed of sound, resulting in a peculiar scaling function. We formulate a discrete boostlet frame using Meyer wavelets and bump functions and examine its sparsity properties. An analysis with experimentally measured fields indicates that discrete boostlet coefficients decay significantly faster and attain superior reconstruction performance than wavelets, curvelets, shearlets, and wave atoms. The results demonstrate that boostlets provide a natural, compact representation system for acoustic waves in space-time.
keywords:
acoustic signal processing , space–time , wave-based modeling , dispersion relation , Poincaré group
@@ 1 Introduction
1.1 From Gabor atoms to Poincaré wavelets
@@
Decades of efforts in applied harmonic analysis have led to remarkable representation systems. In 1946, Gabor presented a theory for time–frequency analysis [3] by applying Heisenberg’s uncertainty principle and establishing an analogy between sound and quanta. Gabor envisioned that a family of parametrized atoms could detect transient components more effectively than Fourier bases. Gabor’s ideas later triggered the development of the wavelet transform in the 1970s, coined initially as the cochlear transform by Zweig [4], and later formalized in various ways by Grossmann, Morlet, Debauchies, and Meyer in the 1980s–1990s [5, 6, 7].
Notable works in the early 1990s by Ali, Antoine, and Gazeau established a framework for square-integrability conditions of unitary representations in space–time. In a two-part study, they introduced reproducing triples and continuous frames [8] and constructed explicit coherent-state wavelet bases on the forward light-cone [9], laying the mathematical foundation for analyzing wave equations via continuous wavelet transforms. Related results were found by Klauer and Streater on cyclic representations of the Poincaré group with a resolution of identity [10], and by Bohnke on the admissibility conditions for Lorentz-group wavelets and tight frames of localized solutions on the light cone [11]. Bernier and Taylor later generalized these ideas for square integrable representations of semidirect products of groups [12]. These works were largely theoretical, but provided a basis for later applications and were eventually consolidated by Ali, Antoine, and Gazeau in a comprehensive book on coherent states and wavelets [13].
By the mid-1990s, Kaiser had introduced the notion of physical wavelets as actual acoustic or electromagnetic pulses rather than abstract basis functions [14, 15]. Kiselev, Perel, and Sidorenko built on these foundations by formulating the Poincaré wavelet transform [16, 17], comprising wavelets that are dilated, Lorentz-boosted, and translated in space–time. In the field of image processing, many other anisotropic transforms were constructed in the late 1990s and early 2000s, e.g., directional wavelets to localize edges [18], ridgelets to localize ridge functions [19], and curvelets [20], shearlets [21], and contourlets [22] to localize curved singularities. In the mid 2000s, research in microlocal analysis showed that curvelets and shearlets resolve the wavefront set [23, 24]. In fact, curvelets [25], shearlets [26], and wave atoms [27] provide optimally sparse representations of the wave propagator in free space. In the late 2000s, Demanet and Ying [27] showed that optimally sparse representations of Fourier integral operators and Green’s function of the wave equation are attained with anisotropic essential support of the order of $2^{-aj}\times 2^{-j/2}$, where $a\in[1/2,1]$. At around the same time, Pinto and Vetterli introduced directional filter banks to approximate wavefronts measured by smoothly-varying 1D microphone arrays [28, 29].
1.2 Historical developments in acoustic simulation and experiments
The analysis and processing of acoustic fields has seen numerous developments since the first computer simulations in the late 1950s [30, 31], when applications of geometrical-acoustics methods, e.g., ray tracing [32], quickly became popular. Some years later, the image source method was developed as an analytical tool to predict acoustical reflections in complex room geometries [33]. In the 1980s, Williams and collaborators made monumental advancements in sound source quantification by introducing acoustic holography [34], an industrial breakthrough for contactless characterization of vibroacoustic sources. By the end of the 1990s, Berkhout et al. studied the analysis [35] and extrapolation [36] of acoustic fields inside rooms in great detail by using analogies from seismic exploration. Numerical approaches such as finite elements [37], finite-difference-time-domain methods [38], and, more recently, spectral element methods [39] and discontinuous Galerkin methods [40], have since been used extensively to approximate the solution by discretizing the computational domain.
Since the emergence of compressed sensing, sparse representations have had impactful applications in acoustic simulations and experiments, including super-resolution, multi-scale analysis, and feature extraction for model-based and machine-learning frameworks. Sampling and reconstructing acoustic fields in space–time is challenging because many data points are required to avoid aliasing. For example, in the human audible range (202020 Hz to 202020 kHz), about 1.61.61.6 million microphone positions are necessary to reconstruct the pressure field over a cubic meter of air. Thus, overcoming sampling limitations has a profound impact on technologies based on the wave equation, such as seismic exploration, medical ultrasound, and vibroacoustic imaging, to mention a few. Notable studies have investigated the sparse reconstruction of acoustic fields with monopole sources [41, 42], plane waves [43], curvelets [44], and shearlets [45]. Finally, data-driven and machine-learning methods, such as physics-informed neural networks [46, 47, 48, 49, 50], deep prior approaches [51, 52, 53], generative models [54, 55, 56, 57], and neural operators [58, 59, 60, 61] have recently achieved impressive results due to their ability to more closely adapt to the structure (spatial or temporal) of the acoustic fields.
@@ 1.3 Contribution of the work
It is worth noting how this study is placed in the existing literature. The mathematical structure of the boostlet transform is similar to wavelet constructions based on the Poincaré group [13]. In particular, boostlets are spatiotemporal atoms parameterized by continuous dilations, Lorentz boosts, and space–time translations—mirroring the parameters of the Poincaré wavelet transform. Despite this similarity, Poincaré-based systems were originally studied in the context of relativistic monochromatic point sources [62] and asymptotic solutions to the wave equation in smoothly inhomogeneous media [63]. In the former context, a Lorentz boost $\theta$ is interpreted from a relativistic perspective as an observer’s velocity, e.g., as in optics, which is inapplicable to acoustic, wave-like signals. In the latter context, a Lorentz boost is interpreted as a characteristic ray that is traced in the medium via $\tan{\beta}=\sinh{\theta}$, which is an asymptotic assumption valid for particle-like (high-frequency) wave propagation. In contrast, the boostlet transform proposed in this work decomposes broadband wavefronts in space–time directly in terms of their phase speed content—determined by passbands of Lorentz boosts—and their frequency content—determined by passbands of space–time scales. Additionally, the boostlet transform takes into consideration finite scales, completing a resolution of identity and thereby ensuring perfect broadband wavefield reconstruction down to zero hertz.
@@
    We apply the principle of analogy from sparsely coding natural images [64, 65] to sparsely coding acoustic waves in space–time. Previous theoretical and experimental results [66] indicate that natural acoustic fields in 2D space–time can be sparsely encoded into band-limited wave shapes supported on frequency-phase speed passbands. This is shown in Figure 1(c), where anti-symmetric hyperbolae localized in phase velocity cones appear to characterize these shapes. This provides experimental evidence supporting that the Poincaré group is a suitable candidate to sparsely represent broadband wavefields in space–time. In particular, this contribution extends to feature-based machine learning in the sense that boostlet features no longer need to be learned during training.
@@
    We introduce the boostlet transform, which decomposes broadband acoustic waves into a collection of band-limited functions on $L^{2}(\mathbb{R}^{2})$, referred to as boostlets, parametrized by the Poincaré group (Lorentz boosts and translations), and isotropic dilations in 2D space–time. The physical meaning of the admissibility condition and the associated scaling function are discussed. A discrete formulation with a Meyer wavelet system is presented. When applied to acoustic fields measured experimentally, discrete boostlet decompositions have a faster coefficient decay compared to benchmark systems such as wavelets, shearlets, and wave atoms.
Figure 1: Illustration of a room-acoustical measurement in 2D space–time. (a) Experimental setup with a sound source emitting a frequency sweep between 100100100 Hz and 4.54.54.5 kHz and a free-field microphone recording the sound pressure. (b) Sound pressure field measured for approximately 8.98.98.9 ms on a line of 100100100 microphone positions spanning 333 m. The zoom-in shows a wave decaying in space–time, entailing a scattered wave. (c) Fourier spectrum of (b). The spectrum shows several hyperbolae localized in a phase velocity cone (dashed lines shown for illustration purposes).
In addition to the abovementioned contributions, the present work aims to improve the accessibility and applicability of space–time wavelets in the context of engineering sciences. While the mathematical foundation of wavelets built from unitary representations of the Poincaré group has been studied since the early 1990s [8, 9, 11, 13, 14, 16, 17, 67], much of this literature is practically inaccessible for researchers working outside harmonic analysis and mathematical physics. In this way, the present work contributes to making space–time wavelet systems more usable and interpretable within the engineering sciences.
@@ 2 Background
Although solutions of the wave equations have a wide frequency range, this study focuses on the acoustic regime between infrasound and ultrasound. Our ambition is that the methodology can be extended to other types of waves and propagation media. This section presents the relevant background on canonical solutions to time-harmonic wave equations, Fourier acoustics theory, encoding dispersion relations with the Poincaré group, and discrete frame theory preliminaries.
@@ 2.1 Wave equation preliminaries
$\Box p(\mathbf{r},t):=\left(\nabla^{2} -\frac{1}{c_{0}^{2}}\frac{\partial^{2}}{\partial t^{2}}\right)p(\mathbf{r},t)=0,$ (1)
@@
$\nabla^{2}=\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}+\frac{\partial^{2}}{\partial z^{2}}.$ (2)
@@
The spatial Fourier transform of Eq. (5) gives the following Dirac delta located at $(\hat{k}_{x},\hat{k}_{y},\hat{k}_{z})$
$\hat{p}(\mathbf{k};\omega_{0})=\int_{\mathbb{R}^{3}}P(\omega_{0})e^{i\,(\mathbf{r}\cdot\hat{\mathbf{k}})}e^{-i(\mathbf{r}\cdot\mathbf{k})}\mathrm{d}\mathbf{r}=P(\omega_{0})\delta(k_{x}-\hat{k}_{x},k_{y}-\hat{k}_{y},k_{z}-\hat{k}_{z}),$ (7)
@@
In the field of room acoustics, experimental evidence [43] suggests that plane wave expansions can sparsely represent stationary acoustic fields at low frequencies (i.e., the modal region). In this frequency region, it is known that the acoustic field in the room is composed of a few modes, which can be represented by a few plane waves. Similar reports have indicated that monopole sources provide sparse representations of compact sound sources [69, 70], which are fully localized in real space and broadband in the wavenumber domain.
@@ 2.2 Acoustic radiation in Fourier space
$\hat{p}(\mathbf{r},\omega_{0})=P(\omega_{0})e^{i(k_{x}x+k_{y}y+k_{z}z)}$ (10)
@@
$\hat{p}(\mathbf{r},\omega_{0})=P(\omega_{0})e^{i(k_{x}x+k_{y}y)}e^{-k_{z}^{\prime}z},$ (11)
@@
For all values of $(k_{x},k_{y})\in\mathbb{R}^{2}$ and a specific frequency $\omega_{0}$, propagating and evanescent waves are thus divided by the so-called radiation circle [68]. As illustrated in Figure 2, the extrusion of the radiation circle as the frequency $\omega\in\mathbb{R}$ varies results in the radiation cone. The presence of this cone suggests that a natural decomposition of the wavefield must preserve this cone.